OpenStudy (anonymous):

Sub. by removing a factor of 1 14bc/b^2-c^2-b-c/b+c

OpenStudy (anonymous):

\[\frac{14bc}{b^2}-\frac{c^2}{-b}-\frac{-c}{b+c}\]Right?

OpenStudy (anonymous):

than denominator is: \[-b^3(b+c)\]

OpenStudy (anonymous):

the problem is that the expression can be interpreted in different ways

OpenStudy (anonymous):

\[\frac{[14bc*(-b(b+c]-[c^2*(b^2(b+c)]-[c*b^2*-b]}{b^2*(-b)*(b+c)}\]

OpenStudy (anonymous):

\[\frac{[14bc*(-b^2-bc)]-[c^2*(b^3+b^2c)]-[-cb^3]}{-b^3(b+c)}\]

OpenStudy (anonymous):

\[\frac{[-14b^3c-14b^2c]-[c^2b^3+b^2c^3]+cb^3}{-b^3(b+c)}\]